OpenStudy (anonymous):

How many grams of magnesium oxide can be produced when 125 g Mg react with 102 g O2? Unbalanced equation: Mg + O2 → MgO Show, or explain, all of your work along with the final answer.

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

M(MgO) = M(Mg) + M(O) = 24,3+ 16 = 40,3 g/mol From the equation : 2 mol of Mg -----> 2 mol of MgO 5,144 mol of Mg -----> 5,144 mol of MgO m(MgO) = M(MgO) x n(MgO) = 40,3 x 5,144 = 207,3 g of MgO

OpenStudy (anonymous):

i dont know if this is right though might want to check @CaseyCarns